# Nonasymptotic Oblivious Relaying and Variable-Length Noisy Lossy Source Coding

### v1

@@ I Introduction
Figure 1: Information bottleneck channel (or oblivious relaying).
@@ II-A Information Bottleneck Channel and Oblivious Relaying
II-B Noisy Lossy Source Coding
The noisy lossy source coding problem, also referred to as remote lossy source coding, was first introduced by Dobrushin and Tsybakov [10]. It was shown to be asymptotically equivalent to the rate-distortion problem with a surrogate distortion between the output of the noisy channel and the decoder’s output, an idea that was further explored in [18]. It reduces to the information bottleneck problem [6] under the logarithmic distortion measure [19]. In [20], the problem was extended to incorporate an $f$-separable distortion measure. Refer to [21] for further extensions with privacy constraints.

### v2

@@ I Introduction
Figure 1: Information bottleneck channel (or oblivious relaying).
@@ II-A Information Bottleneck Channel and Oblivious Relaying
II-B Noiseless and Noisy Lossy Source Coding
In the conventional noiseless lossy source coding setting, the encoder observes the source directly (i.e., $X^{n}=Y^{n}$ in the noisy lossy source coding setting). The optimal length for a fixed-length code was characterized to the second-order in [20, 21], given in terms of the rate-dispersion function [22]. For variable-length codes, the study of $d$-semifaithful codes concerns the setting where the distortion is bounded by $\mathsf{D}$ almost surely [23, 24, 25, 22, 26]. The optimal expected length is $nR(\mathsf{D})+O(\log n)$ [24, 25]. Pointwise bounds on the length have been studied in [22], where the length is shown to be lower-bounded by $nR(\mathsf{D})+\sqrt{n}G_{n}+O(\log n)$ where $G_{n}$ approaches a Gaussian random variable. Variable-length codes allowing a positive probability of excess distortion $\epsilon$ have been studied in [27, 28]. It was shown in [28] that the optimal expected length approaches $n(1-\epsilon)R(\mathsf{D})$ from below.
The noisy lossy source coding problem (where the encoder only has a noisy observation $Y^{n}$ of the source $X^{n}$), also referred to as remote lossy source coding, was first introduced by Dobrushin and Tsybakov [10]. It was shown to be asymptotically equivalent to the rate-distortion problem with a surrogate distortion between the output of the noisy channel and the decoder’s output, an idea that was further explored in [18]. It reduces to the information bottleneck problem [6] under the logarithmic distortion measure [19]. In [29], the problem was extended to incorporate an $f$-separable distortion measure. Refer to [30] for further extensions with privacy constraints.
@@ V Concluding Remarks
VI Acknowledgement
This work was partially supported by two grants from the Research Grants Council of the Hong Kong Special Administrative Region, China [Project No.s: CUHK 24205621 (ECS), CUHK 14209823 (GRF)]. We would like to thank Prof. Ioannis Kontoyiannis for the insightful discussion.